ALL PART OF SINGLE QUESTION

Historical growth rate estimation

Dividend $2.00 $2.50 $2.90 $3.20 $3.60

a. What is the compound growth rate of dividends based on the last five years of dividends data? (2 marks)

Dividend $2.00 $2.50 $2.90 $3.20 $3.60

b. Calculate the year-to-year growth rates in dividends. (2 marks) There are growth periods

c. What is the average year-to-year dividend growth rate? (1 mark)

d. ABC has a retention ratio of 0.75 and a historical return on equity (ROE) of 0.2. Using these two additional pieces of information, calculate an alternative estimate of dividend growth rate, g. (1 mark) Dividend growth model

e. Given that ABC’s share price is currently $65, and the most recent dividend paid is $3.60 per share, use the three growth rates estimated for historical growth (com-pound growth rate, average year-to-year growth rate, and alternative estimate) to calculate the cost of equity using the dividend growth model. (3 marks)

f. Average the three estimated costs of equity in part (e). (1 mark) SML model

g. Given that the firm’s equity beta is 1.6, the risk-free rate is 5%, and the expected return on the market index is 13.5%, calculate its cost of equity using the SML model. (1 mark) WACC calculation

h. Calculate the firm’s average cost of equity by averaging the answers in parts (f) and (g). (1 mark)

i. ABC’s capital structure contains only debt and equity. Given that its debt-equity ratio is 0.8, its cost of debt is 10%, and its marginal tax rate is 35%, calculate the firm’s WACC using the cost of equity calculated in part (h). (2 marks) NPV calculation

j. The firm has a project with an initial cost of $1 million, and annual cash savings of $300,000 for the next five years. The risk adjustment for this project on the WACC is +5%. Calculate the net present value of this project using the WACC calculated above. (4 marks)

Please post answer to e), f) g) h), i) j). thank you!
Answer #1

As you have specifically asked part e to h, the answers of them are provided.

e). Current Price (P0) = $65

Dividend Paid (D0) = $3.6

Growth Rate as computed in part (a), (c), (d) are

- Growth rate as per CAGR (a) = 12.47%

- Growth rate as per Average year-to-year (c) = 15.96%

- Growth rate as per alternate estimate (d) = 15%

Since, you have already computed above growth rate, I'm not providing there solutions.

Dividend growth Model formula :-

Ke = \frac{D_{0}(1+g)}{P_{0}}+g

Using the above formula and different Growth Rate, Calculating Cost of equity:-

- Cost of Equity using Growth Rate (g) as CAGR;-

Ke = \frac{3.6(1+0.1247)}{65}+0.1247

Ke = 18.70%

- Cost of Equity using Growth Rate (g) as Average year-to-year;-

Ke = \frac{3.6(1+0.1596)}{65}+0.1596

Ke = 22.38%

- Cost of Equity using Growth Rate (g) as Alternate Estimate;-

Ke= 3.6(1 + 0.15) 65 - +0.15

Ke = 21.37%

f) Average of the above three Estimate :-

Ke = (18.70+22.38+21.37)/3

Ke = 20.82%

g) Beta of Equity = 1.6

Risk Free Rate (Rf) = 5%

Market return (Rm) = 13.5%

Cost of Equity or expected return using SML Formua:-

Ke = R_{f}+\beta (R_{m}-R_{f})

Ke = 5 +1.6(13.5-5)

Ke = 18.6%

h). Calculating the firm’s average cost of equity by averaging the answers in parts (f) and (g):-

Ke = (20.82+18.6)/2

Ke = 19.71%

i). Debt-Equity Ratio = 0.80

\frac{Debt}{Equity}=\frac{0.80}{1}

So, Debt's is 0.80 & Equity is 1

Cost of Debt = 10%

Tax rate = 35%

Cost of equity (as computed in part(h)) = 19.71%

Computing Firm's WACC:-

WACC=Cost of Equity*(\frac{Equity}{Debt+Equity})+CostofDebt(1-Tax Rate)*(\frac{Debt}{Debt+Equity})

WACC=19.71*(\frac{1}{(1+0.8)})+10(1-0.35)*(\frac{0.8}{(1+0.8)})

WACC = 10.95 + 2.89

WACC = 13.84%

j). Initial cost = $1,000,000

Cash-flows for 5 year each = $300,000

Discount Rate (r) = WACC+5% ;= 18.84%

Calculating the Net Present Value :-

NPV= \frac{CF}{(1+r)^{1}}+\frac{CF}{(1+r)^{2}}+\frac{CF}{(1+r)^{3}}+\frac{CF}{(1+r)^{4}}+\frac{CF}{(1+r)^{5}}-Initial Cost

NPV= \frac{300000}{(1+0.1884)^{1}}+\frac{300000}{(1+0.1884)^{2}}+\frac{300000}{(1+0.1884)^{3}}+\frac{300000}{(1+0.1884)^{4}}+\frac{300000}{(1+0.1884)^{5}}-1000000

NPV = -$79423.3

If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.